# Weak convergence of Bayes estimators under general loss functions

### v1

@@ A.3 Differentiability of the parametric Wasserstein distance
The conditions stated in (Chen and Li,, 2020, Theorem 1) only consider a compact set $\mathcal{X}$ and require that the derivative of the optimal transport map is bounded. These conditions imply the conditions of Theorems 4 and 5. Note that the conditions of Theorems 4 and 5 are just sufficient and not necessary.
Example 7.
Consider the location-scale family of Laplace distributions:
@@
To illustrate the fact that the conditions of Theorems 4 and 5 hold for typical models, we give a concrete example.
Example 8.
@@
on $\mathcal{X}=(1,\infty)$. For any $t,\vartheta\in\Theta$ let $p_{t}$ denote the density of $P_{t}$, $F_{t}$ the c.d.f. of $P_{t}$ , $T_{t}^{\vartheta}$ the optimal transport map, which transports $P_{t}$ to $P_{\vartheta}$ and $\phi_{t}^{\vartheta}$ any solution to the dual problem of the Wasserstein distance. Then, it holds
$p_{t}(x)=\frac{t}{x^{t+1}},\quad\frac{\partial}{\partial t}p_{t}(x)=-\frac{\log(x)t-1}{x^{t+1}},$
$F_{t}(x)=1-x^{-t},\quad\frac{\partial}{\partial t}F_{t}(x)=\frac{\log(x)}{x^{t}},\quad\frac{\partial^{2}}{\partial t^{2}}F_{t}(x)=\frac{\log(x)}{x^{t}}$
$T^{\vartheta}_{t}(x)=F_{\vartheta}^{-1}(F_{t}(x))=x^{t/\vartheta},\quad\frac{\partial}{\partial t}T_{t}^{\vartheta}(x)=\frac{\log(x)x^{t/\vartheta}}{\vartheta},\phi_{t}^{\vartheta}(x)=C+x^{2}-1+\frac{2\vartheta}{t+\vartheta}\left(1-x^{1+t/\vartheta}\right),$
for any $C\in\mathbb{R}$. It can be seen that the conditions of Theorems 4 and 5 are fulfilled, because the logarithm grows slower than $x^{t}$ for any $t>0$ and
@@ Appendix B Technical results
This section includes two technical results, which are consquences of the BvM theorem. The first result shows that the mass of the set $C_{n}^{c}=\left\{h\in H_{n}:\left\lVert h\right\rVert\geq M_{n}\right\}$ with respect to $P_{h_{n}|{\boldsymbol{X}}}$, where $(M_{n})_{n\in\mathbb{N}}$ is any diverging sequence of positive numbers, is asymptotically negligible. As a consequence the posterior distribution asymptotically concentrates on sets of radius of the order $n^{-1/2}$.
@@
$E_{P^{n}_{\vartheta_{0}}}\left[\phi_{n}({\boldsymbol{X}})\right]\to 0\qquad\text{and}\qquad\sup_{\vartheta\in\Theta:\left\lVert\vartheta-\vartheta_{0}\right\rVert\geq\varepsilon}E_{P_{\vartheta}^{n}}\left[1-\phi_{n}({\boldsymbol{X}})\right]\to 0,$
@@
as $n\to\infty$, where $C_{n}:=\{h\in H_{n}:\left\lVert h\right\rVert<M_{n}\}.$
The next result yields under very general conditions the weak convergence of the stochastic process
@@
uniformly over any compact set. It is a consequence of the BvM theorem.
@@
$Z_{n}(t)$ $:=\int_{H_{n}}\ell(t,h)\mathop{}\!\mathrm{d}P_{h_{n}|{\boldsymbol{X}}}(h),$
and
$Z(t)$ $:=\int_{\mathbb{R}^{d}}\ell(t,h)\mathop{}\!\mathrm{d}\mathcal{N}\left(Y,I_{\vartheta_{0}}^{-1}\right)(h),$
@@
The stated two propositions can be proven in the a similar way as Theorem 10.8 from van der Vaart, (1998), but hold for functions even with exponential growth and a proof of them is omitted. They are crucial for the proofs of our asymptotic results, like Propositions 2 and 2.
@@ Appendix C Proofs and technical derivations
In the first case it holds for some fixed $a>0$,
$\ell(t,\vartheta)\geq a\ell(t,\vartheta_{0})-\ell(\vartheta_{0},\vartheta)$
by the triangle inequality of $\rho$ and the fact that for $b_{1},b_{2}\geq 0$ it holds $(b_{1}+b_{2})^{p}\leq 2^{p-1}(b_{1}^{p}+b_{2}^{p})$, if $p\geq 1$ and $(b_{1}+b_{2})^{p}\leq b_{1}^{p}+b_{2}^{p}$, if $0<p<1$. Thus, we choose $a=\min(1,2^{-(p-1)})$. As a consequence,
@@
In the second case, we will show that for any $M>0$ the choice $\delta=M/2$ works. Consider $t,\vartheta\in\Theta$ such that $\left\lVert t-\vartheta_{0}\right\rVert>M$ and $\left\lVert\vartheta-\vartheta_{0}\right\rVert\leq M/2$. We find that
@@
Define $\eta:=\inf_{t,\vartheta:\left\lVert t-\vartheta_{0}\right\rVert>M,\left\lVert\vartheta-\vartheta_{0}\right\rVert\leq\delta}\ell(t,\vartheta)$. By Assumption 2 $\eta$ is strictly positive. We obtain for any sufficiently small $\delta>0$
@@
By Proposition 7 for $g_{n}(h)=1$ and Slutsky’s lemma it holds that
@@ C.2 Proofs for Section 3
We begin the proof by showing that it holds for any $M_{n}\to\infty$ that
@@
By the Assumption (i), we can find $\varepsilon>0$ such that
@@
for all $t,\vartheta\in B_{\varepsilon}(\vartheta_{0})$. In particular, this yields for $\vartheta\in B_{\varepsilon}(\vartheta_{0})$ and any $c_{4}>0,0<c_{5}<2$ that
@@
By assumption (ii) it exists $c_{3}>0,c_{4}>0$ and $0<c_{5}<2$ such that
@@
Then, for $\vartheta\in B_{\varepsilon}(\vartheta_{0})^{C}$ we find that
@@
If we can show that $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ is uniformly tight, then the Theorem follows from the argmax continuous mapping theorem (van der Vaart,, 1998, Corollary 5.58). Let $M_{n}\to\infty$ be any diverging sequence. We will show that for some $r\geq 1$, to be determined later in the proof, the term $\mathbb{P}\left(n^{1/2}\left\lVert\hat{\theta}_{n}-\vartheta_{0}\right\rVert>rM_{n}\right)$ converges to zero, as $n\to\infty$, which implies that $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ is uniformly tight. By assumption, the Bayes estimator is consistent, i.e. $\hat{\theta}_{n}\stackrel{{\scriptstyle P^{n}_{\vartheta_{0}}}}{{\longrightarrow}}\vartheta_{0}$. We find that for any $\varepsilon>0,r\geq 1$
@@
$\tau_{n}:\Theta\to H_{n},\vartheta\mapsto n^{1/2}(\vartheta-\vartheta_{0}).$
For $n\in\mathbb{N}$ such that $rM_{n}n^{-1/2}\leq\varepsilon$ we have for $t,h\in\Omega_{n}$ with $\left\lVert t\right\rVert\leq\varepsilon n^{1/2},\left\lVert h\right\rVert\leq M_{n}$ that $\tau_{n}^{-1}(t),\tau_{n}^{-1}(h)\in B_{\varepsilon}(\vartheta_{0})$. Define $\ell_{n}:H_{n}\times H_{n}\to[0,\infty),(t,h)\mapsto n^{p/2}\ell(\tau_{n}^{-1}(t),\tau_{n}^{-1}(h))$. Consequently, for $t,h\in H_{n}$ such that $\tau_{n}^{-1}(t),\tau_{n}^{-1}(h)\in B_{\varepsilon}(\vartheta_{0})$, we find that
@@
Because $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ minimizes the map $t\mapsto Z_{n}(t)$ almost surely, it holds
@@
Note that by condition (i) and (iii) we can apply Proposition 7 with $g_{n}(h):=\sup_{t\in K}\ell_{n}(t,h)$. Hence, it holds for any $M_{n}\to\infty$
@@
For any $t\in K,h\in C=\{h:\left\lVert h\right\rVert\leq M\}$ we have for some constants $c_{3}>0,c_{4}>0,0<c_{5}<2$
@@
by the extreme value theorem and Assumption (iii). Consequently, it holds
$\left\lVert Z_{n,M}-W_{n,M}\right\rVert_{\infty}\lesssim\left\lVert P_{H_{n}|{\boldsymbol{X}}}-\mathcal{N}(\Delta_{n,\vartheta_{0}},I^{-1}_{\vartheta_{0}})\right\rVert_{\mathsf{TV}}=o_{P^{n}_{\vartheta_{0}}}(1),$
@@
Denote $f_{\mu}=\mathop{}\!\mathrm{d}\mathcal{N}(\mu,I^{-1}_{\vartheta_{0}})/\mathop{}\!\mathrm{d}\lambda^{d}$, where $\mu\in\mathbb{R}^{d}$ and consider an arbitrary sequence $\mu_{n}\to\mu$. It holds for $\left\lVert\cdot\right\rVert_{\infty}$ the sup norm on $\ell^{\infty}(K)$ that
$\left\lVert\mathcal{I}_{n}(\mu_{n})-\mathcal{I}(\mu)\right\rVert_{\infty}$ $\leq\int_{C}\sup_{t\in K}\left\lvert\ell_{n}(t,h)-\ell_{0}(t,h)\right\rvert f_{\mu_{n}}(h)\mathop{}\!\mathrm{d}h+\sup_{t\in K,h\in C}\ell_{0}(t,h)\int_{C}\left\lvert f_{\mu_{n}}(h)-f_{\mu}(h)\right\rvert\mathop{}\!\mathrm{d}h$
@@
This upper bound converges to zero by assumption and the dominated convergence theorem. By the central limit theorem $\Delta_{n,\vartheta_{0}}\stackrel{{\scriptstyle D}}{{\to}}_{P^{n}_{\vartheta_{0}}}Y\sim\mathcal{N}(0,I^{-1}_{\vartheta_{0}})$. As a consequence, the extended continuous mapping theorem (van der Vaart and Wellner,, 2023, Theorem 1.11.1) implies that
$\mathcal{I}_{n}(\Delta_{n,\vartheta_{0}})=W_{n,M}\stackrel{{\scriptstyle D}}{{\to}}_{P^{n}_{\vartheta_{0}}}W_{M}=\mathcal{I}(Y),\quad n\to\infty.$
@@
$Z_{n,M}\stackrel{{\scriptstyle D}}{{\to}}W_{M},\quad n\to\infty$
for any $M>0$. Moreover, we find for $c_{4}>0,0<c_{5}<2$ and $\eta_{2}=\sup_{t\in K}e^{c_{4}\left\lVert t\right\rVert^{c_{5}}}$ and any $y\in\mathbb{R}^{d}$ that by Assumption (iii)
$\sup_{t\in K}\left\lvert Z(t)-W_{M}(t)\right\rvert\lesssim\eta_{2}\int_{h:\left\lVert h\right\rVert>M}e^{c_{4}\left\lVert h\right\rVert^{c_{5}}}\mathop{}\!\mathrm{d}\mathcal{N}(Y,I^{-1}_{\vartheta_{0}})(h).$
@@
Hence,
$W_{M}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}Z,\quad M\to\infty$
in $\ell^{\infty}(K)$. Conclude that there exists $M_{n}\to\infty$ such that $Z_{n,M_{n}}\stackrel{{\scriptstyle D}}{{\to}}Z$ in $\ell^{\infty}(K)$, as $n\to\infty$. Finally, by (12)
$\left\lVert Z_{n}-Z_{n,M_{n}}\right\rVert_{\infty}=\sup_{t\in K}\int_{h:\left\lVert h\right\rVert>M_{n}}\ell_{n}(t,h)\mathop{}\!\mathrm{d}P_{h_{n}|{\boldsymbol{X}}}(h)=o_{P^{n}_{\vartheta_{0}}}(1),$
@@
$Z_{n}\stackrel{{\scriptstyle D}}{{\to}}_{P^{n}_{\vartheta_{0}}}Z$
@@
Define the sequences of loss functions $\ell_{n}(t,h):=n^{p/2}\ell(n^{-1/2}t+\vartheta_{0},n^{-1/2}h+\vartheta_{0})$. We will show that for any compact set $K\subset\mathbb{R}^{d}$ there exists some $0<a<2$ and $\xi>0$ such that for $n$ large enough
$\ell_{n}(t,h)\lesssim e^{\xi\left\lVert t\right\rVert^{a}+\xi\left\lVert h\right\rVert^{a}},\quad t\in K,h\in H_{n}.$
Then, the assertion follows immediately from Lemma 1. By Assumption (i) there exists $p>0$ and $\varepsilon>0$ small enough such that
@@
for all $t,h\in B_{\varepsilon}(\vartheta_{0})$. Moreover, by Assumption (iv) we have for some $\delta>0$ that there exit $c_{2}>0,c_{3}>0,0<c_{4}<2$ such that
@@
Without loss of generality we can assume that $\delta<\varepsilon$, else consider $\delta^{\prime}=\min(\delta,\varepsilon/2)$ instead. Let $N>0$ be large enough such that for $\eta:=\sup_{t\in K}\left\lVert t\right\rVert\leq\delta n^{1/2}/2$ for all $n\geq N$. We separate the range of $h$ between $A_{n}:=\{h\in H_{n}:\sup_{t\in K}\left\lVert t-h\right\rVert\leq\delta n^{1/2}/2\}$ and $B_{n}:=\{h\in H_{n}:\sup_{t\in K}\left\lVert t-h\right\rVert\geq\delta n^{1/2}/2\}$. Note that for any $n\geq N$ and $h\in A_{n}$ we have
$\frac{\left\lVert h\right\rVert}{n^{1/2}}\leq\frac{\left\lVert t\right\rVert}{n^{1/2}}+\frac{\left\lVert t-h\right\rVert}{n^{1/2}}\leq\delta<\varepsilon,\quad t\in K.$
That means $n^{-1/2}h+\vartheta_{0}\in B_{\varepsilon}(\vartheta_{0})$ and $n^{-1/2}t+\vartheta_{0}\in B_{\varepsilon}(\vartheta_{0})$. This yields for any $t\in K$, $h\in A_{n}$ and come constant $c>0$
@@
Now consider any $h\in B_{n}$. We obtain for the $c_{3}>0,0<c_{4}<2$ from (13) and some $\xi_{0},\xi_{1}>0$ the inequality

### v2

@@ A.3 Differentiability of the parametric Wasserstein distance
The conditions stated in (Chen and Li,, 2020, Theorem 1) only consider a compact set $\mathcal{X}$ and require that the derivative of the optimal transport map is bounded. These conditions imply the conditions of Theorems 4 and 5. In addition, note that the conditions of Theorems 4 and 5 are just sufficient and not necessary.
Example 7 (Laplace distribution).
Consider the location-scale family of Laplace distributions
@@
To illustrate the fact that the conditions of Theorems 4 and 5 hold for typical models, we give two examples.
Example 8 (Pareto distribution).
@@
on $\mathcal{X}=(1,\infty)$. For any $t,\vartheta\in\Theta$ let $p_{t}$ denote the density of $P_{t}$, $F_{t}$ the distribution function of $P_{t}$, and $T_{t}^{\vartheta}$ the optimal transport map, which transports $P_{t}$ to $P_{\vartheta}$ and $\Phi(t,\vartheta)$ the set of solutions to the dual formulation of the Wasserstein distance. Then, it holds
$p_{t}(x)=\frac{t}{x^{t+1}},\;\frac{\partial}{\partial t}p_{t}(x)=-\frac{\log(x)t-1}{x^{t+1}},$
$F_{t}(x)=1-x^{-t},\;\frac{\partial}{\partial t}F_{t}(x)=\frac{\log(x)}{x^{t}},\;\frac{\partial^{2}}{\partial t^{2}}F_{t}(x)=\frac{\log(x)}{x^{t}}$
$T^{\vartheta}_{t}(x)=F_{\vartheta}^{-1}(F_{t}(x))=x^{t/\vartheta},\;\frac{\partial}{\partial t}T_{t}^{\vartheta}(x)=\frac{\log(x)x^{t/\vartheta}}{\vartheta},$
${\Phi(t,\vartheta)=\left\{x\mapsto C+x^{2}+\frac{2\vartheta}{t+\vartheta}\left(1-x^{1+t/\vartheta}\right):C\in\mathbb{R}\right\}}.$
It can be seen that the conditions of Theorems 4 and 5 are fulfilled for $\Phi_{0}(t,\vartheta)=\{x\mapsto x^{2}+\frac{2\vartheta}{t+\vartheta}\left(1-x^{1+t/\vartheta}\right)\}$, since the logarithm grows slower than $x^{t}$ for all $t>0$ and
@@
Example 9 (Beta distribution).
Consider the parametric shape family of Beta distributions
$\mathcal{M}=\{P_{t}=\mathrm{Beta}(t,1):t\in\Theta=(0,\infty)\}$
on $\mathcal{X}=(0,1)$. For any $t,\vartheta\in\Theta$, by $p_{t}$ and $F_{t}$ denote the density the distribution function of $P_{t}$, respectively. Let $T_{t}^{\vartheta}$ be the optimal transport map, which transports $P_{t}$ to $P_{\vartheta}$ and $\Phi(t,\vartheta)$ the set of solutions to the dual formulation of the Wasserstein distance. Then, it holds
$p_{t}(x)=tx^{t-1},\;\frac{\partial}{\partial t}p_{t}(x)=x^{t-1}(1+t\log(x)),$
$F_{t}(x)=x^{t},\;\frac{\partial}{\partial t}F_{t}(x)=x^{t}\log(x),\;\frac{\partial^{2}}{\partial t^{2}}F_{t}(x)=x^{t}\log(x)^{2},$
$T^{\vartheta}_{t}(x)=F_{\vartheta}^{-1}(F_{t}(x))=x^{t/\vartheta},\;\frac{\partial}{\partial t}T_{t}^{\vartheta}(x)=\frac{\log(x)x^{t/\vartheta}}{\vartheta},$
$\Phi(t,\vartheta)=\{x\mapsto C+x^{2}-2x^{(t+\vartheta)/\vartheta}:C\in\mathbb{R}\}.$
Thus, with the choice $\Phi_{0}(t,\vartheta)=\{x\mapsto x^{2}-2x^{(t+\vartheta)/\vartheta}\}$, the conditions in Theorems 4 and 5 are fulfilled.
@@ Appendix B Technical results
This supplement includes two technical results, which are consequences of the Bernstein–von Mises (BvM) theorem. The first result shows that the mass of the set $C_{n}^{c}=\left\{h\in H_{n}:\left\lVert h\right\rVert\geq M_{n}\right\}$ with respect to $P_{h_{n}|{\boldsymbol{X}}}$, where $(M_{n})_{n\in\mathbb{N}}$ is any diverging sequence of positive numbers, is asymptotically negligible. As a consequence, the posterior distribution asymptotically concentrates on sets of radius of the order $n^{-1/2}$.
@@
$\mathbb{E}_{P^{n}_{\vartheta_{0}}}\left[\phi_{n}({\boldsymbol{X}})\right]\to 0\qquad\text{and}\qquad\sup_{\vartheta\in\Theta:\left\lVert\vartheta-\vartheta_{0}\right\rVert\geq\varepsilon}\mathbb{E}_{P_{\vartheta}^{n}}\left[1-\phi_{n}({\boldsymbol{X}})\right]\to 0,$
@@
where $C_{n}:=\{h\in H_{n}:\left\lVert h\right\rVert<M_{n}\}.$
The second result yields under very general conditions the weak convergence of the stochastic process
@@
uniformly over any compact set.
@@
$Z_{n}(t):=\int_{H_{n}}\ell(t,h)\mathop{}\!\mathrm{d}P_{h_{n}|{\boldsymbol{X}}}(h),\quad\text{ and }\quad Z(t):=\int_{\mathbb{R}^{d}}\ell(t,h)\mathop{}\!\mathrm{d}\mathcal{N}\left(Y,I_{\vartheta_{0}}^{-1}\right)(h),$
@@
The stated two propositions can be proven in the a similar way as Theorem 10.8 in van der Vaart, (1998), but hold for functions even with exponential growth and a proof of them is thus omitted. They are crucial for the proofs of our asymptotic results, such as Propositions 2 and 2.
@@ Appendix C Proofs and technical derivations
In this supplement, we state the proofs of the results from the main text.
@@
In the first case (i), it holds for $a=\min(1,2^{(1-p)})\in(0,1]$,
$\ell(t,\vartheta_{0})\leq a^{-1}(\ell(t,\vartheta)+\ell(\vartheta_{0},\vartheta))$
by the triangle inequality of $\rho$ and the fact that for $b_{1},b_{2}\geq 0$ and $p>0$, $(b_{1}+b_{2})^{p}\leq 2^{\max\{p-1,0\}}(b_{1}^{p}+b_{2}^{p})$. By rewriting this inequality we find that
$\ell(t,\vartheta)\geq a\ell(t,\vartheta_{0})-\ell(\vartheta_{0},\vartheta).$
As a consequence,
@@
In the second case (ii), we will show that for any $M>0$ the choice $\delta=M/2$ works. Consider $t,\vartheta\in\Theta$ such that $\left\lVert t-\vartheta_{0}\right\rVert>M$ and $\left\lVert\vartheta-\vartheta_{0}\right\rVert\leq M/2$. We obtain by the reverse triangle inequality that
@@
Define $\eta:=\inf_{t,\vartheta:\left\lVert t-\vartheta_{0}\right\rVert>M,\left\lVert\vartheta-\vartheta_{0}\right\rVert\leq\delta}\ell(t,\vartheta)$. By Assumption 2, $\eta$ is strictly positive. We obtain for any sufficiently small $\delta>0$
@@
By Proposition 7, for $g_{n}(h)=1$ and Slutsky’s lemma it holds that
@@ C.2 Proofs for Section 3
We begin the proof by showing that it holds, for any $M_{n}\to\infty$,
@@
By condition (i), we can find $\varepsilon>0$ such that
@@
for all $t,\vartheta\in B_{\varepsilon}(\vartheta_{0})$. In particular, this yields for $\vartheta\in B_{\varepsilon}(\vartheta_{0})$ and $c_{4}>0,0<c_{5}<2$,
@@
By condition (ii) it exists $c_{3}>0,c_{4}>0$ and $0<c_{5}<2$ such that
@@
Then, for $\vartheta\in B_{\varepsilon}(\vartheta_{0})^{C}$, we have
@@
If we can show that $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ is uniformly tight, then the assertion follows from the argmax continuous mapping theorem (van der Vaart,, 1998, Corollary 5.58). Let $M_{n}\to\infty$ be any diverging sequence. We will show that for some $r\geq 1$, to be determined later in the proof, the term $\mathbb{P}\left(n^{1/2}\left\lVert\hat{\theta}_{n}-\vartheta_{0}\right\rVert>rM_{n}\right)$ converges to zero, as $n\to\infty$, which implies that $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ is uniformly tight. By assumption, the Bayes estimator is consistent, i.e. $\hat{\theta}_{n}\stackrel{{\scriptstyle P^{n}_{\vartheta_{0}}}}{{\longrightarrow}}\vartheta_{0}$. We find that for any $\varepsilon>0,r\geq 1$
@@
$\tau_{n}:\Theta\to H_{n},\quad\vartheta\mapsto n^{1/2}(\vartheta-\vartheta_{0}).$
For $n\in\mathbb{N}$ such that $rM_{n}n^{-1/2}\leq\varepsilon$ we have for $t,h\in H_{n}$ with $\left\lVert t\right\rVert\leq\varepsilon n^{1/2},\left\lVert h\right\rVert\leq M_{n}$ that $\tau_{n}^{-1}(t),\tau_{n}^{-1}(h)\in B_{\varepsilon}(\vartheta_{0})$. Define $\ell_{n}:H_{n}\times H_{n}\to[0,\infty),\,(t,h)\mapsto n^{p/2}\ell(\tau_{n}^{-1}(t),\tau_{n}^{-1}(h))$. Consequently, for $t,h\in H_{n}$ such that $\tau_{n}^{-1}(t),\tau_{n}^{-1}(h)\in B_{\varepsilon}(\vartheta_{0})$, we find that
@@
Because $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ minimizes the mapping $t\mapsto Z_{n}(t)$ almost surely, it holds
@@
Note that by conditions (i) and (iii) we can apply Proposition 7 with $g_{n}(h):=\sup_{t\in K}\ell_{n}(t,h)$. Hence, it holds for any $M_{n}\to\infty$
@@
For any $t\in K,h\in C=\{h:\left\lVert h\right\rVert\leq M\}$, we have for some constants $c_{3}>0,c_{4}>0,0<c_{5}<2$
@@
by the extreme value theorem in real analysis and condition (iii). Consequently, it holds
$\left\lVert Z_{n,M}-W_{n,M}\right\rVert_{\infty,K}$
$=\,$ $\sup_{t\in K}\left\lvert\int_{h\in H_{n}:\left\lVert h\right\rVert\leq M}\ell_{n}(t,h)\mathop{}\!\mathrm{d}P_{h_{n}|{\boldsymbol{X}}}(h)-\int_{h\in\mathbb{R}^{d}:\left\lVert h\right\rVert\leq M}\ell_{n}(t,h)\mathop{}\!\mathrm{d}\mathcal{N}(\Delta_{n,\vartheta_{0}},I^{-1}_{\vartheta_{0}})(h)\right\rvert$
$\lesssim\,$ $\left\lvert P_{h_{n}|{\boldsymbol{X}}}(\{h\in\mathbb{R}^{d}:\left\lVert h\right\rVert\leq M\})-\mathcal{N}(\Delta_{n,\vartheta_{0}},I^{-1}_{\vartheta_{0}})(\{h\in\mathbb{R}^{d}:\left\lVert h\right\rVert\leq M\})\right\rvert$
$\lesssim\,$ $\left\lVert P_{H_{n}|{\boldsymbol{X}}}-\mathcal{N}(\Delta_{n,\vartheta_{0}},I^{-1}_{\vartheta_{0}})\right\rVert_{\mathsf{TV}}=o_{P^{n}_{\vartheta_{0}}}(1),$
@@
Denote $f_{\mu}=\mathop{}\!\mathrm{d}\mathcal{N}(\mu,I^{-1}_{\vartheta_{0}})/\mathop{}\!\mathrm{d}\lambda^{d}$, where $\mu\in\mathbb{R}^{d}$ and consider an arbitrary sequence $\mu_{n}\to\mu$. It holds for $\left\lVert\cdot\right\rVert_{\infty,K}$ the sup norm on $\ell^{\infty}(K)$ that
$\left\lVert\mathcal{I}_{n}(\mu_{n})-\mathcal{I}(\mu)\right\rVert_{\infty,K}$
$\leq$ $\int_{C}\sup_{t\in K}\left\lvert\ell_{n}(t,h)-\ell_{0}(t,h)\right\rvert f_{\mu_{n}}(h)\mathop{}\!\mathrm{d}h+\sup_{t\in K,h\in C}\ell_{0}(t,h)\int_{C}\left\lvert f_{\mu_{n}}(h)-f_{\mu}(h)\right\rvert\mathop{}\!\mathrm{d}h$
@@
This upper bound converges to zero by condition (ii) and the dominated convergence theorem. By the central limit theorem $\Delta_{n,\vartheta_{0}}\stackrel{{\scriptstyle D}}{{\to}}_{P^{n}_{\vartheta_{0}}}Y\sim\mathcal{N}(0,I^{-1}_{\vartheta_{0}})$. As a consequence, the extended continuous mapping theorem (van der Vaart and Wellner,, 2023, Theorem 1.11.1) implies that
$\mathcal{I}_{n}(\Delta_{n,\vartheta_{0}})=W_{n,M}\stackrel{{\scriptstyle D}}{{\longrightarrow}}_{P^{n}_{\vartheta_{0}}}W_{M}=\mathcal{I}(Y),\quad n\to\infty.$
@@
$Z_{n,M}\stackrel{{\scriptstyle D}}{{\longrightarrow}}W_{M},\quad n\to\infty$
for any $M>0$. Moreover, we find for $c_{4}>0,0<c_{5}<2$ and $\eta_{2}=\sup_{t\in K}e^{c_{4}\left\lVert t\right\rVert^{c_{5}}}$ and any $y\in\mathbb{R}^{d}$ that by condition (iii)
$\sup_{t\in K}\left\lvert Z(t)-W_{M}(t)\right\rvert$ ${=\sup_{t\in K}\int_{h:\left\lVert h\right\rVert>M}\ell_{0}(t,h)\mathop{}\!\mathrm{d}\mathcal{N}(Y,I^{-1}_{\vartheta_{0}})(h)}$
$\lesssim\eta_{2}\int_{h:\left\lVert h\right\rVert>M}e^{c_{4}\left\lVert h\right\rVert^{c_{5}}}\mathop{}\!\mathrm{d}\mathcal{N}(Y,I^{-1}_{\vartheta_{0}})(h).$
@@
Hence, $W_{M}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}Z,\,M\to\infty$ in $\ell^{\infty}(K)$. Conclude that there exists $M_{n}\to\infty$ such that $Z_{n,M_{n}}\stackrel{{\scriptstyle D}}{{\to}}Z$ in $\ell^{\infty}(K)$, as $n\to\infty$. Finally, by (15)
$\left\lVert Z_{n}-Z_{n,M_{n}}\right\rVert_{\infty,K}=\sup_{t\in K}\int_{h:\left\lVert h\right\rVert>M_{n}}\ell_{n}(t,h)\mathop{}\!\mathrm{d}P_{h_{n}|{\boldsymbol{X}}}(h)=o_{P^{n}_{\vartheta_{0}}}(1),$
@@
$Z_{n}\stackrel{{\scriptstyle D}}{{\longrightarrow}}_{P^{n}_{\vartheta_{0}}}Z$
@@
We define the sequences of loss functions
$\ell_{n}(t,h):=n^{p/2}\ell(n^{-1/2}t+\vartheta_{0},n^{-1/2}h+\vartheta_{0}),$
and will employ Lemma 1. To this end, note that requirement (ii) in Lemma 1 follows directly from condition (iii) here. Due to compactness of $K$, there exists $N>0$ such that $t/n^{1/2}+\vartheta\in B_{\delta}(\vartheta_{0})$ for all $t\in K,n\geq N$. By condition (ii), we have
$\int_{\Theta}\sup_{t\in K}\ell_{n}\!\bigl(t,n^{1/2}(\vartheta-\vartheta_{0})\bigr)\,\mathrm{d}\Pi(\vartheta)\leq n^{p/2}\int_{\Theta}\sup_{t\in B_{\delta}(\vartheta_{0})}\ell(t,\vartheta)\,\mathrm{d}\Pi(\vartheta),\quad n\geq N.$
This shows requirement (i) in Lemma 1.
We will next show that for any compact set $K\subset\mathbb{R}^{d}$ there exist some $0<a<2$ and $\xi>0$ such that for $n$ large enough
$\ell_{n}(t,h)\lesssim e^{\xi\left\lVert t\right\rVert^{a}+\xi\left\lVert h\right\rVert^{a}},\quad t\in K,h\in H_{n},$
i.e. requirement (iii) in Lemma 1. By condition (i), there exist $p>0$ and some sufficiently small $\varepsilon>0$ such that
@@
for all $t,h\in B_{\varepsilon}(\vartheta_{0})$. Moreover, by condition (iv), we have for some $\delta>0$ that there exist $c_{2}>0,c_{3}>0,0<c_{4}<2$ such that
@@
Without loss of generality, we can assume that $\delta<\varepsilon$; otherwise consider $\delta^{\prime}=\min(\delta,\varepsilon/2)$ instead. Let $N>0$ be large enough such that for $\eta:=\sup_{t\in K}\left\lVert t\right\rVert\leq\delta n^{1/2}/2$ for all $n\geq N$. We separate the range of $h$ between $A_{n}:=\{h\in H_{n}:\sup_{t\in K}\left\lVert t-h\right\rVert\leq\delta n^{1/2}/2\}$ and $B_{n}:=\{h\in H_{n}:\sup_{t\in K}\left\lVert t-h\right\rVert\geq\delta n^{1/2}/2\}$. For any $n\geq N$ and $h\in A_{n}$, we have
$\frac{\left\lVert h\right\rVert}{n^{1/2}}\leq\frac{\left\lVert t\right\rVert}{n^{1/2}}+\frac{\left\lVert t-h\right\rVert}{n^{1/2}}\leq\delta<\varepsilon,\quad t\in K,$
i.e. $n^{-1/2}h+\vartheta_{0}\in B_{\varepsilon}(\vartheta_{0})$ and $n^{-1/2}t+\vartheta_{0}\in B_{\varepsilon}(\vartheta_{0})$. This yields, for any $t\in K$, $h\in A_{n}$ and come constant $c>0$,
@@
Now consider any $h\in B_{n}$. We obtain with $c_{3}>0,0<c_{4}<2$ in (16) and some $\xi_{0},\xi_{1}>0$,